# P. Guiraud, M. Hernández, A. Meyroneinc and A. Nogueira

### v1

@@ 3.3 Monotonicity properties of $\phi$
$<1-\lambda-d+\frac{1-\lambda}{\lambda}\left((1-\lambda)\sum_{k=1}^{\infty}\lambda^{k}(k-1)-(1-\lambda)\sum_{k=k_{1}+1}^{\infty}\lambda^{k}+(1-\lambda)\sum_{k=k_{1}+1}^{\infty}\lambda^{k}\right)$
$<1-\lambda-d+\frac{1-\lambda}{\lambda}\left((1-\lambda)\lambda\sum_{k=1}^{\infty}\lambda^{k-1}k-(1-\lambda)\sum_{k=1}^{\infty}\lambda^{k}\right)$
$<1-\lambda-d+\frac{1-\lambda}{\lambda}\left(\frac{\lambda}{1-\lambda}-\lambda\right)$
$<1-d.$
@@
Let $\rho\in[0,1]$, $\alpha\in[0,1]$ and $\delta\in(1-\lambda-d,1-\lambda)$. Let $\eta$ be defined as in (46).
@@
1)1)1 ) Suppose $\alpha\neq 1-\rho$ and let $y\in[0,1)$. As $\delta\leq\delta(\rho,\alpha)$, we have $\phi((1-\rho)^{-})\leq\eta$ with strict inequality for $\delta\neq\delta(\rho,\alpha)$. Therefore, it follows from Proposition 3.6 that $\phi(y)\leq\eta$ if $y<1-\rho$ and the inequality is strict if $\delta\neq\delta(\rho,\alpha)$ or $\rho\in(0,1)\setminus\mathbb{Q}$. On the other hand, as $\delta\geq\delta(\rho^{-},\alpha)$, we have $\phi(1-\rho)\geq\eta$ and $\phi(y)\geq\eta$ for $y\geq 1-\rho$. The proof of the case $\alpha=1-\rho$ is analogous.
2)2)2 ) Suppose $\alpha\neq 1-\rho$. If $\phi(y)<\eta$ for every $y\in[0,1-\rho)$, then $\phi((1-\rho)^{-})\leq\eta$. If $\rho\in(0,1)\cap\mathbb{Q}$, then there exists $y\in[0,1-\rho)$ such that $\phi(y)=\phi((1-\rho)^{-})$, by Lemma 3.8. It follows that $\phi((1-\rho)^{-})\leq\eta$ and that $\phi((1-\rho)^{-})<\eta$ if $\rho\in(0,1)\cap\mathbb{Q}$. We deduce that $\delta\leq\delta(\rho,\alpha)$ with $\delta\neq\delta(\rho,\alpha)$ if $\rho\in(0,1)\cap\mathbb{Q}$. On the other hand, if $\phi(y)\geq\eta$ for every $y\in[1-\rho,1)$, then $\phi((1-\rho))\geq\eta$ and $\delta\geq\delta(\rho^{-},\alpha)$. The proof of the case $\alpha=1-\rho$ is analogous. ∎
@@
    $[\phi(\alpha^{-}),\phi(\alpha)]\subset(0,\eta_{1})\quad\text{if}\quad\rho\notin\mathbb{Q}\quad\text{and}\quad[\phi(\alpha^{-}),\phi(\alpha)]\subset[0,\eta_{1}]\quad\text{if}\quad\rho\in\mathbb{Q}.$ (48)
@@
    $[\phi(\alpha^{-}),\phi(\alpha)]\subset(\eta_{2},1)\quad\text{if}\quad\rho\notin\mathbb{Q}\quad\text{and}\quad[\phi(\alpha^{-}),\phi(\alpha)]\subset[\eta_{2},1]\quad\text{if}\quad\rho\in\mathbb{Q}.$ (49)
@@
Therefore, we always have $[\phi(\alpha^{-}),\phi(\alpha)]\cap[\eta_{1},\eta_{2}]\neq\emptyset$, since $\phi(\alpha^{-})\leq\phi(\alpha)$ by monotonicity of $\phi$. Also,
@@ 6.1 Proof of the equivalences
Now, we start the proof. In the following 1)1)1 ), 2)2)2 ) and 3)3)3 ) refer to the items of Theorem 2.9.
1)2)1$\Rightarrow$ 2)1 ) ⇒ 2 ). If (20) holds, then by (a) and (c) of Remark 6.1, we have
@@
which proves (21) and item 2)2)2 ).
2)3).2$\Rightarrow$ 3).2 ) ⇒ 3 ) . Recall that for any $\alpha\in[0,1]$ we have
@@
see Lemma 3.2. As $\phi([0,1))\subset[0,1)$ when 2)2)2 ) holds, we have that $\{\phi(y)\}=\phi(\{y\})$ and $\left\lfloor\phi(y)\right\rfloor=\left\lfloor y\right\rfloor$, therefore (21) implies that
@@
3)1).3$\Rightarrow$ 1).3 ) ⇒ 1 ) . By hypothesis we have
@@
Applying Corollary 5.4 we obtain that $\phi([0,1))\subset[0,1)$ as desired. We have proved that 3)1)3$\Rightarrow$ 1)3 ) ⇒ 1 ) for $(\delta,a)\in M_{1}$. The same can be shown with an analogous proof for $(\delta,a)\in M_{3}$. Finally, if $(\delta,a)\in M_{2}$, then $\alpha=1-\rho$ and
@@
since $a\in[\eta_{1},\eta_{2}]$. From Corollary 5.4 we deduce that $\phi([0,1))\subset[0,1)$ and we conclude that 3)1)3$\Rightarrow$ 1)3 ) ⇒ 1 ) also for $(\delta,a)\in M_{2}$.
@@ 6.2 Image of $[0,1)$ by $\phi$
Let $\delta\in[0,1]$, $\rho\in(0,1)\setminus\mathbb{Q}$, $\alpha\in[0,1]$ and denote $\phi=\phi_{\delta,\rho,\alpha}$. Then, the closure of $\phi([0,1))$ is a Cantor set. Moreover, if $\{\alpha+k\rho\}\neq 0$ for every $k\geq 1$ and $\delta$ satisfy (20), then the Cantor set extremities are the points 00 and $1$.
@@
It remains to show that $\overline{\phi([0,1))}$ is totally disconnected. Let $A\subset\overline{\phi([0,1))}$ be a connected set and suppose that $a:=\inf A<b:=\sup A$. Since $A$ is connected $A\neq\{a,b\}$. Thus, there exists $x\in(a,b)$ and $y\in[0,1)$ such that $\phi(y)=x$. As the sequence ($(\{l\rho\})_{l\in\mathbb{N}}$ is dense in $[0,1)$ and $\phi$ is right continuous at $y$, there exists $l\geq 1$, such that $\phi(\{l\rho\})\in(a,b)$. Thus,
@@
Also,
$\phi(\{l\rho\})-\phi(\{l\rho\}^{-})=\frac{1}{\lambda}\sum_{k=1}^{\infty}\lambda^{k}((1-\lambda-d)\mathbb{1}_{\mathbb{Z}}(\{l\rho\}-k\rho)+d\mathbb{1}_{\mathbb{Z}+\alpha}(\{l\rho\}-k\rho))>0.$
Therefore, $(\phi(\{l\rho\}),\phi(\{l\rho\}^{-}))$ is a gap of $\phi$ and
$A=(A\cap[a,\phi(\{l\rho\})])\cup(A\cap[\phi(\{l\rho\}^{-}),b])$
@@ 6.2.1 Image of $\phi$ and attractor of $f$ for $\rho$ irrational
Then, $D^{-}_{\alpha,\rho}$ and $D^{+}_{\alpha,\rho}$ contain at most one element and $D^{-}_{\alpha,\rho}=\emptyset$ or $D^{+}_{\alpha,\rho}=\emptyset$. Moreover,
@@
    $[f(a^{-}),f(a))=[\phi_{\alpha}(\{\rho+\alpha\}^{-})),\phi_{\alpha}(\{\rho+\alpha\})\quad\text{if}\quad\alpha\neq 1-\rho$ (56)
@@
    $[0,f(a))\cup[f(a^{-}),1)=[0,\phi(0))\cup[\phi(1^{-}),1)\quad\text{if}\quad\alpha=1-\rho.$ (57)
@@
    If $D^{+}_{\alpha,\rho}=\emptyset$, then $\phi(0^{-})=\phi(0)=0$ and $[f(1^{-}),f(0))=[\phi_{\alpha}(\rho^{-}),\phi_{\alpha}(\rho))$.
@@
As $\rho\in(0,1)\setminus\mathbb{Q}$, the set $D^{-}_{\rho,\alpha}\cup D^{+}_{\rho,\alpha}=\{k\in\mathbb{Z}^{*}:\{k\rho+\alpha\}=0\}$ contains at most one element. Therefore, $D^{-}_{\rho,\alpha}=\emptyset$ if $D^{+}_{\rho,\alpha}\neq\emptyset$ and $D^{+}_{\rho,\alpha}=\emptyset$ if $D^{-}_{\rho,\alpha}\neq\emptyset$.
@@
$\eta=\eta_{1}$ if $\alpha\in[0,1-\rho)$, $\eta=\eta_{2}$ if $\alpha\in(1-\rho,1]$ and $\eta=a$ if $\alpha=1-\rho$. As $\phi$ is increasing and $\{k\rho+\beta\}\notin\{\alpha,1-\rho\}$, the interval $[\phi_{\alpha}(\{k\rho+\beta\}^{-}),\phi_{\alpha}(\{k\rho+\beta\})$ does not contain $\eta$ or $a$. Therefore, $f$ is increasing and continuous in this interval. We deduce that
@@
Now, as $\{k\rho+\beta\}\neq 0$, there exits a positive sequence $(\epsilon_{n})_{n\in\mathbb{N}}$ converging to zero and such that $10\leq\{k\rho+\beta\}-\epsilon_{n}<$ 1 for all $n\in\mathbb{N}$. Then, using the continuity of $f$ and (21), we obtain
@@
3) If $D_{\alpha,\rho}^{-}=\emptyset$, then $k\rho\notin\mathbb{Z}+\alpha$ for every $k\geq 1$. It follows from Lemma 4.3 and (20) that $\phi(\alpha^{-})=\phi(\alpha)=a$. Also, by Proposition 2.5, we have $\alpha\in(0,1)$ if and only if $a\in(0,1)$, since $\rho\in(0,1)\setminus\mathbb{Q}$. Now, suppose $a\in(0,1)$ and let $(\epsilon_{n})_{n\in\mathbb{N}}$ be a positive sequence converging to zero and such that $10\leq\alpha-\epsilon_{n}<$ 1 for all $n\in\mathbb{N}$. Then, $\phi(\alpha-\epsilon_{n})\in[0,a)$ for every $n\in\mathbb{N}$ and $\phi(\alpha-\epsilon_{n})\to a$. Therefore,
@@
4) If $D_{\alpha,\rho}^{+}=\emptyset$, then $-k\rho\notin\mathbb{Z}+\alpha$ for every $k\geq 1$. It follows from Lemma 4.2 that $\delta(\rho^{-},\alpha)=\delta(\rho,\alpha)$ which implies that $\phi(0^{-})=\phi(0)$ and $\phi(0)=0$ by (20). Now, let $(\epsilon_{n})_{n\in\mathbb{N}}$ be a positive sequence converging to zero and such that $10\leq 1-\epsilon_{n}<$ 1 for all $n\in\mathbb{N}$. Then, $\phi(1-\epsilon_{n})\in[0,1)$ for every $n\in\mathbb{N}$ and $\phi(1-\epsilon_{n})\to 1$, since $\phi(1^{-})=\phi(0^{-})+1=\phi(1)$. Therefore,
@@
Case 1: $D^{-}_{\rho,\alpha}=D^{+}_{\rho,\alpha}=\emptyset$. Then, $\phi(0^{-})=\phi(0)=0$ and $\phi(1^{-})=\phi(0^{-})+1=1$ by 4). It follows that $H_{0}=\emptyset$. Also, by 4) and 1) we have
@@
If $a\in\{0,1\}$, then $\alpha\in\{0,1\}$ and $H_{2}^{k}=H_{1}^{k}$ for every $k\geq 1$. Thus, there is nothing more to prove. If $a\in(0,1)$, then $\alpha\in(0,1)$. Also, $D^{+}_{\rho,\alpha}=\emptyset$ implies $\alpha\neq 1-\rho$. It follows from 3) and 2) that
@@
Case 2: $D^{-}_{\rho,\alpha}\neq\emptyset$. Then, $D^{+}_{\rho,\alpha}=\emptyset$ and there exists a unique $l\geq 1$ such that $\{\alpha-l\rho\}=0$. This implies $\alpha\notin\{0,1-\rho,1\}$ and $\{k\rho\}=\{(k-l)\rho+\alpha\}$ for every $k\geq 1$. So, we have
@@
Now, as $D^{+}_{\rho,\alpha}=\emptyset$, we have $\phi(0^{-})=\phi(0)=0$ and $\phi(1^{-})=\phi(0^{-})+1=1$, by 4). It follows that $H_{0}=\emptyset$. On the other hand, as $\phi(\alpha^{-})\leq a\leq\phi(\alpha)$, we have
@@
and $H_{2}^{1}=f(H_{1}^{l})\cup G_{2}^{0}$. On the other hand, item 2) ensures that $H_{2}^{m+1}=f(H_{2}^{m})$ for all $m\geq 1$, since $D^{+}_{\rho,\alpha}=\emptyset$. So, we obtain
@@
Still because $D^{+}_{\rho,\alpha}=\emptyset$, item 4) gives

### v3

@@ 3.3 Monotonicity properties of $\phi$
$=1-\lambda-d+\frac{1-\lambda}{\lambda}\left((1-\lambda)\sum_{k=1}^{\infty}\lambda^{k}(k-1)-(1-\lambda)\sum_{k=k_{1}+1}^{\infty}\lambda^{k}+(1-\lambda)\sum_{k=k_{1}+1}^{\infty}\lambda^{k}\right)$
$=1-\lambda-d+\frac{1-\lambda}{\lambda}\left((1-\lambda)\lambda\sum_{k=1}^{\infty}\lambda^{k-1}k-(1-\lambda)\sum_{k=1}^{\infty}\lambda^{k}\right)$
$=1-\lambda-d+\frac{1-\lambda}{\lambda}\left(\frac{\lambda}{1-\lambda}-\lambda\right)$
$=1-d.$
@@
Let $\rho\in[0,1]$, $\alpha\in[0,1]$ and $\delta\in(1-\lambda-d,1)$. Let $\eta$ be defined as in (50).
@@
$1)$ Suppose $\alpha\neq 1-\rho$ and let $y\in[0,1)$. As $\delta\leq\delta(\rho,\alpha)$, we have $\phi((1-\rho)^{-})\leq\eta$ with strict inequality for $\delta\neq\delta(\rho,\alpha)$. Therefore, it follows from Proposition 3.6 that $\phi(y)\leq\eta$ if $y<1-\rho$ and the inequality is strict if $\delta\neq\delta(\rho,\alpha)$ or $\rho\in(0,1)\setminus\mathbb{Q}$. On the other hand, as $\delta\geq\delta(\rho^{-},\alpha)$, we have $\phi(1-\rho)\geq\eta$ and $\phi(y)\geq\eta$ for $y\geq 1-\rho$. The proof of the case $\alpha=1-\rho$ is analogous.
$2)$ Suppose $\alpha\neq 1-\rho$. If $\phi(y)<\eta$ for every $y\in[0,1-\rho)$, then $\phi((1-\rho)^{-})\leq\eta$. If $\rho\in(0,1)\cap\mathbb{Q}$, then there exists $y\in[0,1-\rho)$ such that $\phi(y)=\phi((1-\rho)^{-})$, by Lemma 3.8. It follows that $\phi((1-\rho)^{-})\leq\eta$ and that $\phi((1-\rho)^{-})<\eta$ if $\rho\in(0,1)\cap\mathbb{Q}$. We deduce that $\delta\leq\delta(\rho,\alpha)$ with $\delta\neq\delta(\rho,\alpha)$ if $\rho\in(0,1)\cap\mathbb{Q}$. On the other hand, if $\phi(y)\geq\eta$ for every $y\in[1-\rho,1)$, then $\phi((1-\rho))\geq\eta$ and $\delta\geq\delta(\rho^{-},\alpha)$. The proof of the case $\alpha=1-\rho$ is analogous. ∎
@@
    $\phi(\alpha^{-}),\phi(\alpha)]\subset(0,\eta_{1})\quad\text{if}\quad\rho\notin\mathbb{Q}\quad\text{and}\quad[\phi(\alpha^{-}),\phi(\alpha)]\subset[0,\eta_{1}]\quad\text{if}\quad\rho\in\mathbb{Q}.$ (52)
@@
    $\phi(\alpha^{-}),\phi(\alpha)]\subset(\eta_{2},1)\quad\text{if}\quad\rho\notin\mathbb{Q}\quad\text{and}\quad[\phi(\alpha^{-}),\phi(\alpha)]\subset[\eta_{2},1]\quad\text{if}\quad\rho\in\mathbb{Q}.$ (53)
@@
Therefore, we always have $\phi(\alpha^{-}),\phi(\alpha)]\cap[\eta_{1},\eta_{2}]\neq\emptyset$, since $\phi(\alpha^{-})\leq\phi(\alpha)$ by monotonicity of $\phi$. Also,
@@ 6.1 Proof of the equivalences
Now, we start the proof. In the following $1)$, $2)$ and $3)$ refer to the items of Theorem 2.10.
$1)\Rightarrow 2)$. If (24) holds, then by (a) and (c) of Remark 6.1, we have
@@
which proves (25) and item $2)$.
$2)\Rightarrow 3).$ Recall that for any $\alpha\in[0,1]$ we have
@@
see Lemma 3.2. As $\phi([0,1))\subset[0,1)$ when $2)$ holds, we have that $\{\phi(y)\}=\phi(\{y\})$ and $\left\lfloor\phi(y)\right\rfloor=\left\lfloor y\right\rfloor$, therefore (25) implies that
@@
$3)\Rightarrow 1).$ By hypothesis we have
@@
Applying Corollary 5.4 we obtain that $\phi([0,1))\subset[0,1)$ as desired. We have proved that $3)\Rightarrow 1)$ for $(\delta,a)\in M_{1}$. The same can be shown with an analogous proof for $(\delta,a)\in M_{3}$. Finally, if $(\delta,a)\in M_{2}$, then $\alpha=1-\rho$ and
@@
since $a\in[\eta_{1},\eta_{2}]$. From Corollary 5.4 we deduce that $\phi([0,1))\subset[0,1)$ and we conclude that $3)\Rightarrow 1)$ also for $(\delta,a)\in M_{2}$.
@@ 6.2 Image of $[0,1)$ by $\phi$
Let $\delta\in[0,1]$, $\rho\in(0,1)\setminus\mathbb{Q}$, $\alpha\in[0,1]$ and denote $\phi=\phi_{\delta,\rho,\alpha}$. Then, the closure of $\phi([0,1))$ is a Cantor set. Moreover, if $\{\alpha+k\rho\}\neq 0$ for every $k\geq 1$ and $\delta$ satisfy (24), then the Cantor set extremities are the points 0 and $1$.
@@
It remains to show that $\overline{\phi([0,1))}$ is totally disconnected. Let $A\subset\overline{\phi([0,1))}$ be a non-empty connected set and suppose that $a:=\inf A<b:=\sup A$. Since $A$ is connected $A\neq\{a,b\}$. Thus, there exists $x\in(a,b)\subset A$ and $y\in[0,1)$ such that $\phi(y)=x$. As the sequence $(\{l\rho\})_{l\in\mathbb{N}}$ is dense in $[0,1)$ and $\phi$ is right continuous and increasing, there exists $l\geq 1$ such that $y<\{l\rho\}$ and $\phi(\{l\rho\}^{-}),\phi(\{l\rho\})\in(a,b)$. Thus,
@@
Also, by (35)
$\phi(\{l\rho\})-\phi(\{l\rho\}^{-})=\frac{1}{\lambda}\sum_{k=1}^{\infty}\lambda^{k}((1-\lambda-d)\mathbb{1}_{\mathbb{Z}}(\{l\rho\}-k\rho)+d\mathbb{1}_{\mathbb{Z}+\alpha}(\{l\rho\}-k\rho))>0,$
since $\{l\rho\}-k\rho\in\mathbb{Z}$ for $k=l$. Therefore, $(\phi(\{l\rho\}^{-}),\phi(\{l\rho\}))$ is a gap of $\phi$ and
$A=(A\cap[a,\phi(\{l\rho\}^{-})])\cup(A\cap[\phi(\{l\rho\}),b])$
@@ 6.2.1 Image of $\phi$ and attractor of $f$ for $\rho$ irrational
Then, $D^{-}_{\alpha,\rho}$ and $D^{+}_{\alpha,\rho}$ contain at most one element and $D^{-}_{\alpha,\rho}=\emptyset$ or $D^{+}_{\alpha,\rho}$$\emptyset$. Moreover,
@@
    $f(a^{-}),f(a))=[\phi_{\alpha}(\{\rho+\alpha\}^{-})),\phi_{\alpha}(\{\rho+\alpha\})\quad\text{if}\quad\alpha\neq 1-\rho$ (60)
@@
    $0,f(a))\cup[f(a^{-}),1)=[0,\phi(0))\cup[\phi(1^{-}),1)\quad\text{if}\quad\alpha=1-\rho.$ (61)
@@
    If $D^{+}_{\alpha,\rho}$$\emptyset$, then $\phi(0^{-})=\phi(0)=0$ and $f(1^{-}),f(0))=[\phi_{\alpha}(\rho^{-}),\phi_{\alpha}(\rho))$.
@@
As $\rho\in(0,1)\setminus\mathbb{Q}$, the set $D^{-}_{\rho,\alpha}\cup D^{+}_{\rho,\alpha}=\{k\in\mathbb{Z}^{*}:\{k\rho+\alpha\}=0\}$ contains at most one element. Therefore, $D^{-}_{\rho,\alpha}=\emptyset$ if $D^{+}_{\rho,\alpha}\neq\emptyset$ and $D^{+}_{\rho,\alpha}$$\emptyset$ if $D^{-}_{\rho,\alpha}\neq\emptyset$.
@@
$\eta=\eta_{1}$ if $\alpha\in[0,1-\rho)$, $\eta=\eta_{2}$ if $\alpha\in(1-\rho,1]$ and $\eta=a$ if $\alpha=1-\rho$. As $\phi$ is increasing and $\{k\rho+\beta\}\notin\{\alpha,1-\rho\}$, the interval $\phi_{\alpha}(\{k\rho+\beta\}^{-}),\phi_{\alpha}(\{k\rho+\beta\})$ does not contain $\eta$ or $a$. Therefore, $f$ is increasing and continuous in this interval. We deduce that
@@
Now, as $\{k\rho+\beta\}\neq 0$, there exists a positive sequence $(\epsilon_{n})_{n\in\mathbb{N}}$ converging to zero and such that $0\leq\{k\rho+\beta\}-\epsilon_{n}<1$ for all $n\in\mathbb{N}$. Then, using the continuity of $f$ and (25), we obtain
@@
3) If $D_{\alpha,\rho}^{-}=\emptyset$, then $k\rho\notin\mathbb{Z}+\alpha$ for every $k\geq 1$. It follows from Lemma 4.3 and (24) that $\phi(\alpha^{-})=\phi(\alpha)=a$. Also, by Proposition 2.4, we have $\alpha\in(0,1)$ if and only if $a\in(0,1)$, since $\rho\in(0,1)\setminus\mathbb{Q}$. Now, suppose $a\in(0,1)$ and let $(\epsilon_{n})_{n\in\mathbb{N}}$ be a positive sequence converging to zero and such that $0\leq\alpha-\epsilon_{n}<1$ for all $n\in\mathbb{N}$. Then, $\phi(\alpha-\epsilon_{n})\in[0,a)$ for every $n\in\mathbb{N}$ and $\phi(\alpha-\epsilon_{n})\to a$. Therefore,
@@
4) If $D_{\alpha,\rho}^{+}=\emptyset$, then $k\rho\notin\mathbb{Z}+\alpha$ for every $k\geq 1$. It follows from Lemma 4.2 that $\delta(\rho^{-},\alpha)=\delta(\rho,\alpha)$ which implies that $\phi(0^{-})=\phi(0)$ and $\phi(0)=0$ by (24). Now, let $(\epsilon_{n})_{n\in\mathbb{N}}$ be a positive sequence converging to zero and such that $0\leq 1-\epsilon_{n}<1$ for all $n\in\mathbb{N}$. Then, $\phi(1-\epsilon_{n})\in[0,1)$ for every $n\in\mathbb{N}$ and $\phi(1-\epsilon_{n})\to 1$, since $\phi(1^{-})=\phi(0^{-})+1=\phi(1)$. Therefore,
@@
Case 1: $D^{-}_{\rho,\alpha}=D^{+}_{\rho,\alpha}$$\emptyset$. Then, $\phi(0^{-})=\phi(0)=0$ and $\phi(1^{-})=\phi(0^{-})+1=1$ by 4). It follows that $H_{0}=\emptyset$. Also, by 4) and 1) we have
@@
If $a\in\{0,1\}$, then $\alpha\in\{0,1\}$ and $H_{2}^{k}=H_{1}^{k}$ for every $k\geq 1$. Thus, there is nothing more to prove. If $a\in(0,1)$, then $\alpha\in(0,1)$. Also, $D^{+}_{\rho,\alpha}$$\emptyset$ implies $\alpha\neq 1-\rho$. It follows from 3) and 2) that
@@
Case 2: $D^{-}_{\rho,\alpha}\neq\emptyset$. Then, $D^{+}_{\rho,\alpha}$$\emptyset$ and there exists a unique $l\geq 1$ such that $\{\alpha-l\rho\}=0$. This implies $\alpha\notin\{0,1-\rho,1\}$ and $\{k\rho\}=\{(k-l)\rho+\alpha\}$ for every $k\geq 1$. So, we have
@@
Now, as $D^{+}_{\rho,\alpha}$$\emptyset$, we have $\phi(0^{-})=\phi(0)=0$ and $\phi(1^{-})=\phi(0^{-})+1=1$, by 4). It follows that $H_{0}=\emptyset$. On the other hand, as $\phi(\alpha^{-})\leq a\leq\phi(\alpha)$, we have
@@
and $H_{2}^{1}=f(H_{1}^{l})\cup G_{2}^{0}$. On the other hand, item 2) ensures that $H_{2}^{m+1}=f(H_{2}^{m})$ for all $m\geq 1$, since $D^{+}_{\rho,\alpha}$$\emptyset$. So, we obtain
@@
Still because $D^{+}_{\rho,\alpha}$$\emptyset$, item 4) gives